Suppose a government report about a countrys employment situation reveals the following data (all values are in millions of people) 260Working-age population 90Employed people 28People not working but actively looking for work 34 Unemployed people who have stopped looking for work 108 People not working and not seeking work If the country in question is the United States, these data are obtained from From the data above, it can be determined that this countrys labor force is million people. It can be determined that this countrys unemployment rate is | %. (Round your response to one decimal place.) The data indicate that the countrys participation rate is % Round your response to one decimal place.

The options for the first question are
1. Household survey's

2. private-sector payroll's

3. unemployment office registrations
Answer #1

Unemployment office registration

Data on unemployment in United States is taken from unemployment office registration.

Country's labor force= 90+28= 118

unemployment rate= 28*100/118= 23.7%

participation rate= 118*100/260= 45.4%